# Optimization of foreground moment deprojection for semi-blind CMB polarization reconstruction

### v1

@@ 4.2 Optimization of moment deprojection
Table 2 presents the optimal moment configurations determined by ocMILC across different sky regions and angular scales for each foreground data set. For each needlet band and GNILC cluster (sky region with uniform foreground complexity), the table includes information on the optimal number and combination of moments for deprojection, along with the corresponding optimal values of deprojection coefficients and dust pivot parameters. Examining these optimal configurations allows us to draw conclusions consistent with the discussions in previous sections:
@@
5 Discussion
The results in Section 4 demonstrated the capability of ocMILC to handle diverse dust models efficiently, thanks to the optimization of moment deprojection. This was achieved by integrating information from the GNILC diagnosis and optimizing the selection of moments, pivot values and deprojection coefficients across various sky regions and angular scales. In this section, our focus shifts to assessing the pipeline’s robustness against varied models of Galactic synchrotron emission, while maintaining a fixed thermal dust emission model (d1) to isolate the impact of synchrotron model variations. Additionally, we investigate the pipeline’s response to potential polarization of the anomalous microwave emission (AME).
@@ 5 Discussion
$B$ modes with different synchrotron models
Model $m_{\textrm{fgds}}$ ($f_{\textrm{sky}}(\%)$)
$j=1$ $j=2$ $j=3$ $j=4$ $j=5$
d1s1 $6$ $4\ (13),5\ (87)$ $3\ (4),4\ (95),5\ (1)$ $2\ (11.5),3\ (28),4\ (60.5)$ $1\ (0.7),2\ (68),3\ (16.3),4\ (15)$
d1s5 $6$ $5$ $3\ (4),4\ (52),5\ (44)$ $2\ (1),3\ (38),4\ (50),5\ (11)$ $1\ (0.7),2\ (58.5),3\ (25.5),4\ (15.3)$
d1s7 $6$ $5$ $3\ (4),4\ (52),5\ (44)$ $2\ (1),3\ (38),4\ (50),5\ (11)$ $1\ (0.7),2\ (59),3\ (25),4\ (15.3)$
$B$ modes with AME with different polarization fractions
Model $m_{\textrm{fgds}}$ ($f_{\textrm{sky}}(\%)$)
$j=1$ $j=2$ $j=3$ $j=4$ $j=5$
$\Pi_{\textrm{AME}}=0.5\%$ $7$ $4\ (6.4),5\ (64),6\ (29.6)$ $3\ (4),4\ (89.5),5\ (6.5)$ $2\ (11),3\ (28.5),4\ (60.5)$ $1\ (3),2\ (67),$ $3\ (25.5),4\ (4.5)$
$\Pi_{\textrm{AME}}=1\%$ $7$ $5\ (27.5),6\ (72.5)$ $3\ (4),4\ (74),$ $5\ (10.5),6\ (11.5)$ $2\ (10),3\ (29.6),4\ (55.2),$ $5\ (5),6\ (0.2)$ $2\ (71.5),3\ (27),4\ (1.5)$
$\Pi_{\textrm{AME}}=2\%$ $8$ $6\ (81.5),7\ (18.5)$ $3\ (4),4\ (45),5\ (24),6\ (27)$ $1\ (0.8),2\ (62.8),3\ (19.6)$ $4\ (14.8),5\ (2)$ $1\ (2),2\ (51),$ $3\ (20.4),4\ (26.6)$
@@
After evaluating the effects of more complex synchrotron and AME models in the GNILC diagnosis, we proceed to assess their impact on NILC and ocMILC CMB $B$-mode reconstruction in Figure 8. This is to determine whether there is a need to extend the optimization procedure of ocMILC to accommodate the higher complexity of synchrotron and AME polarized emissions. Indeed, despite ocMILC’s optimization including synchrotron moments, the current pipeline maintains a fixed pivot value of $\bar{\beta}_{\textrm{s}}=-3$ for the synchrotron spectral index (Section 4.2). However, a higher variability of this parameter across the sky may necessitate optimizing the corresponding pivot values, similar to what was done for dust spectral parameters. Furthermore, the framework of moment expansion has not yet been extended to AME [26], so the current ocMILC pipeline does not attempt to deproject any AME moments, leaving them for blind variance minimization.
@@ 6 Conclusions
Deprojecting moments of Galactic emission with the semi-blind cMILC method theoretically allows for minimizing foreground variance rather than overall variance. However, the associated noise penalty, including the increased variance of unconstrained foreground moments, does not always guarantee optimal outcomes for the most complex foreground scenarios. Specifically, deprojecting foreground moments in the natural order of the Taylor expansion may not be optimal, particularly for the moments that dominate the overall variance, which may be better suited for blind variance minimization. To address this trade-off, we introduced an optimized variant of cMILC, named ocMILC (Section 4), which refines the deprojection process through a data-driven approach by optimizing the number and selection of moments to deproject, as well as pivot values and deprojection coefficients.
In ocMILC, the optimization of the number of moments to deproject, depending on sky areas and angular scales, relies on the GNILC diagnosis of data complexity (Section 3), which maps out the number of independent observable modes of Galactic emission, $m_{\textrm{fgds}}$, in the given data set across the sky and at different angular scales (Figure 2 and Table 1). This diagnosis informs ocMILC on the optimal number of moments to deproject locally before paying significant noise penalties. This strategy is corroborated by the results obtained in Figure 5. The specific subset of moments to deproject locally is also optimized to retain some dominant moments for blind variance minimization, while selectively deprojecting higher-order moments that are not predominantly captured by covariance estimation. The optimization of local pivot values of the dust spectral parameters across different sky regions and ranges of angular scales allows for accommodating local moment expansions associated with local averaging processes. Finally, allowing the deprojection coefficients $\epsilon_{k}$ to be non-zero for partial deprojection of certain moments offers additional flexibility for limiting noise penalties. In addition, it allows to better handle with foreground modes not properly described by the moment expansion of a single MBB SED. The optimal pivot values, subset of moments, and deprojection coefficients are determined to minimize the denoised output variance (Equation 4.12), and hence the residual foreground variance, in each GNILC cluster.
@@ Acknowledgments
Appendix A Results for PICO CBE sensitivities
Throughout this paper, we adopted the baseline instrumental configuration of PICO, consistent with required specifications detailed in the PICO report [6], which provides an overall sensitivity of $0.87$ $\mu K\cdot\text{arcmin}$. In this appendix, we present NILC and ocMILC outcomes for the PICO d1 data set, employing the alternative PICO configuration known as CBE (Current Best Estimate) [6, 30], featuring an enhanced overall sensitivity of $0.61$ $\mu K\cdot\text{arcmin}$. The results are shown in Figure 9, showcasing the angular power spectra of residuals averaged over all realizations and corresponding posteriors of the estimated tensor-to-scalar ratio. Even with the increased sensitivity of the PICO CBE configuration, ocMILC still outperforms NILC, exhibiting lower foreground residuals across all angular scales with only a moderate rise in noise contamination. As anticipated, the noise residuals in both NILC and ocMILC CMB $B$-mode reconstructions are lower compared to those obtained with the baseline PICO configuration (displayed in the top panel of Figure 6). Consequently, this leads to lower upper bounds on the estimated tensor-to-scalar ratio ($r<0.00025$ at $95\%$ CL), as shown in the right panel of Figure 9. Furthermore, ocMILC improves upon the biased upper bound obtained with NILC for such a large sky fraction ($70\%$). The ocMILC constraint aligns with the quoted forecast in [30] for PICO CBE sensitivities, albeit for a significantly larger sky fraction compared to [30].

### v2

@@ 4.2.2 CMB $B$-mode reconstruction
Table 2 presents the optimal moment configurations determined by ocMILC across different sky regions and angular scales for each foreground data set. For each needlet band and GNILC cluster (sky region with uniform foreground complexity), the table includes information on the optimal number and combination of moments for deprojection, along with the corresponding optimal values of deprojection coefficients and dust pivot parameters. Examining these configurations allows us to draw conclusions consistent with the discussions in previous sections:
@@
Throughout this paper, we adopted the baseline instrumental configuration of PICO, consistent with required specifications detailed in the PICO report [6], which provides an overall sensitivity of $0.87$ $\mu K\cdot\text{arcmin}$. To complement our results, we present NILC and ocMILC outcomes for the PICO d1 data set, employing the alternative PICO configuration known as CBE (Current Best Estimate) [6, 30], featuring an enhanced overall sensitivity of $0.61$ $\mu K\cdot\text{arcmin}$. The results are shown in Figure 8, showcasing the angular power spectra of residuals averaged over all realizations and corresponding posteriors of the estimated tensor-to-scalar ratio. Even with the increased sensitivity of the PICO CBE configuration, ocMILC still outperforms NILC, exhibiting lower foreground residuals across all angular scales with only a moderate rise in noise contamination. As anticipated, the noise residuals in both NILC and ocMILC CMB $B$-mode reconstructions are lower compared to those obtained with the baseline PICO configuration (displayed in the top panel of Figure 6). Consequently, this leads to lower upper bounds on the estimated tensor-to-scalar ratio ($r<0.00025$ at $95\%$ CL), as shown in the right panel of Figure 8. Furthermore, ocMILC improves upon the biased upper bound obtained with NILC for such a large sky fraction ($70\%$). The ocMILC constraint aligns with the quoted forecast in [30] for PICO CBE sensitivities, albeit for a significantly larger sky fraction compared to [30].
4.3 Sensitivity to AME and synchrotron emission
The results in Section 4.2.2 demonstrated the capability of ocMILC to handle diverse dust models efficiently, thanks to the optimization of moment deprojection. This was achieved by integrating information from the GNILC diagnosis and optimizing the selection of moments, pivot values and deprojection coefficients across various sky regions and angular scales. In this section, our focus shifts to assessing the pipeline’s robustness against varied models of Galactic synchrotron emission, while maintaining a fixed thermal dust emission model (d1) to isolate the impact of synchrotron model variations. Additionally, we investigate the pipeline’s response to potential polarization of the anomalous microwave emission (AME).
@@ 4.3 Sensitivity to AME and synchrotron emission
$B$-modes with different synchrotron models
{adjustwidth}
-1.5cm             Model $m_{\textrm{fgds}}$ ($f_{\textrm{sky}}(\%)$) $j=1$ $j=2$ $j=3$ $j=4$ $j=5$ d1s1 $6$ $4\ (13),5\ (87)$ $3\ (4),4\ (95),5\ (1)$ $2\ (11.5),3\ (28),4\ (60.5)$ $1\ (0.7),2\ (68),3\ (16.3),4\ (15)$ d1s5 $6$ $5$ $3\ (4),4\ (52),5\ (44)$ $2\ (1),3\ (38),4\ (50),5\ (11)$ $1\ (0.7),2\ (58.5),3\ (25.5),4\ (15.3)$ d1s7 $6$ $5$ $3\ (4),4\ (52),5\ (44)$ $2\ (1),3\ (38),4\ (50),5\ (11)$ $1\ (0.7),2\ (59),3\ (25),4\ (15.3)$
$B$-modes with AME with different polarization fractions
{adjustwidth}-2.7cm                      Model $m_{\textrm{fgds}}$ ($f_{\textrm{sky}}(\%)$) $j=1$ $j=2$ $j=3$ $j=4$ $j=5$ $\Pi_{\textrm{AME}}=0.5\%$ $7$ $4\ (6.4),5\ (64),6\ (29.6)$ $3\ (4),4\ (89.5),5\ (6.5)$ $2\ (11),3\ (28.5),4\ (60.5)$ $1\ (3),2\ (67),$ $3\ (25.5),4\ (4.5)$ $\Pi_{\textrm{AME}}=1\%$ $7$ $5\ (27.5),6\ (72.5)$ $3\ (4),4\ (74),$ $5\ (10.5),6\ (11.5)$ $2\ (10),3\ (29.6),4\ (55.2),$ $5\ (5),6\ (0.2)$ $2\ (71.5),3\ (27),4\ (1.5)$ $\Pi_{\textrm{AME}}=2\%$ $8$ $6\ (81.5),7\ (18.5)$ $3\ (4),4\ (45),5\ (24),6\ (27)$ $1\ (0.8),2\ (62.8),3\ (19.6)$ $4\ (14.8),5\ (2)$ $1\ (2),2\ (51),$ $3\ (20.4),4\ (26.6)$
@@
After evaluating the effects of more complex synchrotron and AME models in the GNILC diagnosis, we proceed to assess their impact on NILC and ocMILC CMB $B$-mode reconstruction in Figure 9. This is to determine whether there is a need to extend the optimization procedure of ocMILC to accommodate the higher complexity of synchrotron and AME polarized emissions. Indeed, despite ocMILC’s optimization including synchrotron moments, the current pipeline maintains a fixed pivot value of $\bar{\beta}_{\textrm{s}}=-3$ for the synchrotron spectral index (Section 3.2.2). However, a higher variability of this parameter across the sky may necessitate optimizing the corresponding pivot values, similar to what was done for dust spectral parameters. Furthermore, the framework of moment expansion has not yet been extended to AME [26], so the current ocMILC pipeline does not attempt to deproject any AME moments, leaving them for blind variance minimization.
@@ 5 Conclusions
Deprojecting moments of Galactic emission with the semi-blind cMILC method theoretically allows for minimizing foreground variance rather than overall variance. However, the associated noise penalty, including the increased variance of unconstrained foreground moments, does not always guarantee optimal outcomes for the most complex foreground scenarios. Specifically, deprojecting foreground moments in the natural order of the Taylor expansion may not be optimal, particularly for the moments that dominate the overall variance, which may be better suited for blind variance minimization. To address this trade-off, we introduced an optimized variant of cMILC, named ocMILC (Section 3.2), which refines the deprojection process through a data-driven approach by optimizing the number and selection of moments to deproject, as well as pivot values and deprojection coefficients.
In ocMILC, the optimization of the number of moments to deproject, depending on sky areas and angular scales, relies on the GNILC diagnosis of data complexity (Section 3.2.1), which maps out the number of independent observable modes of Galactic emission, $m_{\textrm{fgds}}$, in the given data set across the sky and at different angular scales (Figure 3 and Table 1). This diagnosis informs ocMILC on the optimal number of moments to deproject locally before paying significant noise penalties. This strategy is corroborated by the results obtained in Figure 5. The specific subset of moments to deproject locally is also optimized to retain some dominant moments for blind variance minimization, while selectively deprojecting higher-order moments that are not predominantly captured by covariance estimation. The optimization of local pivot values of the dust spectral parameters across different sky regions and ranges of angular scales allows for accommodating local moment expansions associated with local averaging processes. Finally, allowing the deprojection coefficients $\epsilon_{k}$ to be non-zero for partial deprojection of certain moments offers additional flexibility for limiting noise penalties. In addition, it allows to better handle foreground modes not properly described by the moment expansion of a single MBB SED. The optimal pivot values, subset of moments, and deprojection coefficients are determined to minimize the denoised output variance (Equation 3.24), and hence the residual foreground variance, in each GNILC cluster.